### v2

# ORCID: 0009-0007-8992-9682

Abstract
Large language models have gained significant traction and popularity in recent times, extending their usage to code-generation tasks. While this field has garnered considerable attention, the exploration of testing and evaluating the robustness of code generation models remains an ongoing endeavor. Previous studies have primarily focused on code generation models specifically for the Python language, overlooking other widely-used programming languages. In this research, we conduct a comprehensive comparative analysis to assess the robustness performance of several prominent code generation models. Furthermore, we investigate how their performance varies across different programming languages. To accomplish this, we introduce perturbations in four key areas of the prompt: DocString, function name, syntax, and format. We have compiled and released a dedicated dataset for this purpose. This work presents our experimental findings, shedding light on the performance of code generation models in various scenarios.
Keywords:
Large Language Model Code Generation Robustness Testing
1 Introduction
In the last few years, Large language models (LLMs) have shown remarkable performance in a wide range of natural language processing (NLP) tasks, including machine translation, text generation, and summarization. Inspired by the success of LLMs in NLP, researchers recently proposed to apply LLMs to solve code-related tasks, such as generating code from a natural language (NL) prompt and completing code from a program context (e.g., code and NL comment).
Despite these advancements, the robustness of the LLMs on code generation tasks has not been extensively explored. LLMs are known to suffer from the issue of robustness in NLP tasks, i.e., studies(Samanta and Mehta, 2017; Gao et al., 2018; Li et al., 2018; Jin et al., 2020; Li et al., 2020) in NLP find that a slight change in the input can cause the model to generate a different prediction from the original result. Meanwhile, some recent studies  (Wang et al., 2023; Mastropaolo et al., 2023) focusing on LLM for code generation find that the Large Language Code generation model may also be sensitive to small perturbations on the prompt. For instance, Figure 1 demonstrates that even a minor function name modification can cause the model to produce incorrect code, although the non-perturbed prompt produced code that passed the test cases. Therefore, to test the robustness of code generation models, Wang et al. (2023) propose the first evaluation benchmark, ReCode, by customizing more than 30 semantic-preserved perturbations on the prompts (e.g., docstrings, function names), and then they evaluate the robustness of three code generation models (i.e., CodeGen, InCoder, and GPT-J).
However, the robustness evaluation (Wang et al., 2023) was limited to the Python language, raising the concern of the generalization to other programming languages, which is important to evaluate and show different facets (Austin et al., 2021), especially considering the diverse language features and different training resources across different programming languages. On the other hand, the evaluation benchmark they used is inadequate regarding the test cases, which may potentially impact the findings (Liu et al., 2023). For example, in Figure 2, the generated code in Java passes against the test suite in HumanEval-X but fails against the stronger test suite in EvalPlus (Liu et al., 2023).
Figure 1: A modification in the function name causes incorrect code generation by CodeGen-6B-Multi model
Figure 2: A generated code by CodeGen-6B-Multi model which passes the test cases of HumanEval-X but fails on the test cases from EvalPlus.
In this paper, we expand the current robustness assessment and revisit the findings from prior research to understand whether they still hold with different programming languages and under more rigorous testing. We aim to provide more insights to guide future research on improving the robustness of LLM-powered code generation. Specifically, we extend ReCode (Wang et al., 2022) from two aspects: First, in addition to Python, we evaluated the LLMs on more programming languages, including Java, C++, and JavaScript. Second, knowing that the test cases in HumanEval are inadequate, we construct an enhanced multi-lingual dataset of test cases based on EvalPlus (Liu et al., 2023). EvalPlus (Liu et al., 2023), an enhancement of the HumanEval (Chen et al., 2021) benchmark, introduces additional corner cases and test cases, leading to a more rigorous evaluation that expands a set of challenges, resulting in a higher failure rate for generated code by Large Language Models. Since EvalPlus test cases and solutions are in Python, we convert them to Java, C++, and JavaScript and manually update the solutions accordingly. The source code and benchmark dataset are publicly available. Our primary contributions are as follows.
  1. (1)
    Our work involves the expansion of ReCode, a robustness testing framework initially developed to evaluate the performance of code generation models using Python datasets. In our research, we extend the capabilities of ReCode to encompass additional programming languages such as Java, C++, and JavaScript.
  2. (2)
    We conduct a comparative analysis to evaluate the performance of code generation models when exposed to perturbed datasets in Java, C++, and JavaScript languages. To the best of our knowledge, no previous research has conducted a comparative study of this nature, evaluating code generation models across multiple languages using perturbed datasets.
  3. (3)
    We release the robustness evaluation datasets for Java, C++, and JavaScript languages, which are perturbed based on HumanEval-X, providing valuable resources that can be utilized for future research and related endeavors.
  4. (4)
    Our study reveals prevalent error types and vulnerabilities in code generation models when subjected to perturbations, highlighting the need for language-specific considerations and improvements in model design and evaluation.
  5. (5)
    We undertake the translation of the EvalPlus dataset from Python to Java, C++, and JavaScript. During this process, we also curate the test cases to ensure their compatibility with the mentioned programming languages. Additionally, we update the canonical solutions within the existing HumanEval-X datasets to align them with the translated test cases. The result of our efforts is a finalized dataset, encompassing translations and curated test cases, which we make publicly available.
In this study, we expand upon the work of ReCode by including Java, C++, and JavaScript languages, allowing for a comparative analysis. We assess the robustness of code generation models across these three languages, employing extended perturbations specific to each language. Four code generation models are experimented upon using nominal and perturbed datasets, evaluating their reactions and measuring robustness. We identify effective perturbation types and explore potential differences in model robustness based on programming languages. Additionally, we analyze prevalent error types resulting from the perturbations.
Artifact. We share the data and code in this replication package [^1] https://github.com/frabbisw/robustextended.
2 Background and Related Work
In this section, we discuss the necessary background information to enhance the clarity of our work. And then we present the recent work that is related to our research.
2.1 Background
In this part, we introduce two fundamental concepts essential to this work: (1) code generation, which is the task of generating source code based on the given natural language (NL) description and other information, (2) robustness testing, which aims to measure the ability of a system to function correctly despite the perturbation of the input.
Code generation aims to generate a code snippet based on a given prompt, which usually contains information about the method’s signature, an NL description of the code’s functionality, and test cases. The model is required to generate a complete code snippet based on the given information. Then the quality of the generated source code is evaluated against the unit tests. For example, in Figure 3, the prompt contains a docstring describing the method’s functionality and a method signature. The rest of the lines in that example are generated code by a code generation model.
Figure 3: An example of a generated code by CodeGen-6B-Multi snippet from a sample prompt
Robustness testing assesses the ability of a model to function correctly despite the perturbation of the input. During testing, a negligible perturbation is performed on the original prompt, and then the modified prompt is fed into the model to generate a new prediction, which is later compared with the original prediction.
In this paper, we mainly focus on the robustness of the code generation model, and during the testing, the prompt is modified slightly (e.g., by changing one letter from lowercase to uppercase), and then the model is asked to generate new code based on the new prompt. The correctness of the newly generated code is examined against the same test cases.
Formally, let $\mathcal{F\left(\cdot\right)}$ be the code generation model, that takes in a prompt $\mathbf{p}=\left(p_{1},p_{2},\dots,p_{N}\right)$, and generates a complete function, $\mathbf{y}=\left(y_{1},y_{2},\dots,y_{M}\right)$, of which the correctness is tested against the set of $K$ test cases $T=\{t_{1},t_{2},...,t_{K}\}$. To test the robustness of the model $\mathcal{F\left(\cdot\right)}$, a new prompt $\mathbf{p}_{adv}$ is generated by slightly perturbing the original prompt (i.e., $\mathbf{p}$), which should follow the following requirements:
$Sim(\mathbf{p}_{adv},\mathbf{p})\geq\epsilon$ (1)
Where $Sim(.)\rightarrow(0,1)$ is the similarity between the original and the perturbed prompt. To preserve the semantic meaning of the original prompt, this similarity value must be larger than the threshold. And the robustness of the model $\mathcal{F\left(\cdot\right)}$ is measured by executing the newly generated function $\mathcal{F}\left(\mathbf{p}_{adv}\right)$ against the same set of test cases $T$.
Code generation models arrive from Large Language Models (Radford et al., 2019)(Brown et al., 2020) using few-shot learning for the code-generation tasks. These language models generally complete a sentence or verse. For code generation tasks, a prompt is pushed into a language model where the prompt has the information about a code’s parameters, test cases, signature, and other documentation, however, the implementation is missing. The language model completes the code by implementing it using the information from the given prompt.
Robustness testing assesses the ability of a component or system to function correctly despite invalid inputs or stressful conditions. It evaluates how well a system handles unexpected scenarios and aims to identify failures caused by errors or unusual inputs. The goal is to ensure the system remains stable, avoids crashes, and maintains its functionality. Robustness testing enhances software reliability, reduces the risk of failure, and optimizes performance in abnormal circumstances. Generally, two types of robustness testing are common in Machine Learning, which are as follows:
Black-Box Testing: Black box testing involves limited access to a model’s inputs and the dataset for labels or confidence intervals. It examines the external behavior of software or machine learning models, emphasizing functionality, user interface usability, and compliance with external standards. Operating without knowledge of internal workings, this testing method validates the desired output. To assess the robustness of code generation models, black box testing is employed. This ensures the models can handle unexpected scenarios and typographical errors, bolstering their reliability and performance.
White-Box Testing: White box robustness testing offers testers extensive access to a model’s output, architecture, gradients, and training data, facilitating a thorough examination of code, algorithms, and data structures. While our research primarily focuses on black box testing, it is important to acknowledge that white box testing provides valuable insights into the system’s inner workings and remains essential for evaluating software and machine learning models. This methodology enables the identification of vulnerabilities and weaknesses, but it requires expertise due to its in-depth analysis of the system’s internals.
Our exploration of related works is organized into four distinct sections, each providing a comprehensive overview of the following areas.
2.2 Related Work
Researchers have proposed various approaches for the task of code generation, and recently, the LLM-based code generation models have achieved impressive performance.
Nijkamp et al. (2022) introduces CodeGen, a series of transformer-based code generation models with various parameters for multiple languages, where the 16 billion parameter variant performs the best. CodeGeex by Zheng et al. (2023) is another multi-language-based series of models for code generation, where the largest model consists of 13 billion parameters. Other notable works in the field of code generation include Codex by Chen et al. (2021), Incoder by Fried et al. (2022), and GPT-J by Wang and Komatsuzaki (2021). Additionally, large language models such as GPT-3 by Brown et al. (2020) have shown potential for code-generation tasks. Furthermore, works such as CodeBERT by Feng et al. (2020) and PLBART by Ahmad et al. (2021) focus on training large language models specifically on code.
Meanwhile, to evaluate the performance of code generation models, prior work has proposed several benchmarks. The HumanEval dataset by Chen et al. (2021) consists of handcrafted prompts and test cases designed for Python, making it suitable for evaluating generated Python code. An expansion of this dataset is HumanEval-X(Zheng et al., 2023), which retains the same prompts but includes test cases for multiple programming languages. Additionally, Evalplus by Liu et al. (2023), an enhanced version of Humaneval, featuring a broader range of test cases and updated canonical solutions was proposed to effectively identify incorrect code, especially in challenging corner cases.
To our knowledge, the first work of generating adversarial examples was done by Szegedy et al. (2013), where they apply a specific barely noticeable perturbation that is discovered by maximizing the network’s prediction error, causing the image classifier network to incorrectly classify an image. Some other similar works (Goodfellow et al., 2014) (Moosavi-Dezfooli et al., 2016) (Kurakin et al., 2018) (Papernot et al., 2016a) (Chib and Greenberg, 1995) have been found to generate adversarial examples with different approaches.
Our research has uncovered various attacks on Natural Language Models that are relevant to our work.  Papernot et al. (2016b) pioneer the crafting of Adversarial Input Sequences for Recurrent Neural Networks in Natural Language Processing. Similarly, Samanta and Mehta (2017) propose an approach involving modifications to original text samples, such as the deletion or replacement of important words and the introduction of new words. Gao et al. (2018) develop a black-box attack technique that alters characters to generate small perturbations, deceiving deep-learning classifiers. Li et al. (2018) focus on character-level and word-level attacks, considering word importance ranking and employing perturbations on NLP classifier models. Jin et al. (2020) introduce TextFooler, a semantic black-box attack challenging the robustness of the BERT model.
Additionally, BERT-ATTACK by Li et al. (2020) targets vulnerable words in input sequences, generating substitutes using BERT. Iyyer et al. (2018) design SCPN, a strategy for sentence paraphrasing without compromising quality. Naik et al. (2018) propose an evaluation methodology using artificially generated stress tests to assess inferential judgments. In software engineering, Zhou et al. (2022) develop ACCENT, testing the robustness of code comment generation by replacing identifiers while preserving functionality and proposing a new training method.
These studies collectively contribute to the understanding of attacks in Natural Language Processing and software engineering domains, advancing techniques for generating adversarial samples and evaluating model robustness.
Although many attacks have been proposed to challenge the robustness of the Machine Learning Models, few are concerned with the code generation model.
ReCode, introduced by Wang et al. (2022), stands as a comprehensive testing framework designed to evaluate the robustness of code generation models specifically tailored for Python datasets. In our study, we closely align with and build upon the foundations established by ReCode. This framework puts forth a diverse set of 28 perturbations encompassing function names, docstrings, and code. The approach outlined in ReCode is evaluated against various popular code generation models, providing valuable insights into their performance and vulnerabilities. By following the methodology outlined in ReCode, we aim to extend the understanding of code generation model robustness in our research.
An additional notable study we encountered is CodeAttack by Jha and Reddy (2022), which presents a black-box attack methodology specifically tailored to exploit the vulnerabilities of Programming language models through code-specific adversarial attacks. This approach leverages the inherent structure of code to generate adversarial code samples that are both efficient and effective while remaining undetectable. The authors identify vulnerable tokens by masking words and inputting them into the models. CodeAttack primarily emphasizes three areas: code-to-code translation, code-to-code repair, and code-to-comment generation, highlighting the diverse applications of their approach within the realm of programming.
In their recent work, Zhuo et al. (2023) conduct an empirical study focusing on the prompt-based adversarial robustness of CODEX. The authors uncover the vulnerability of the model to meticulously crafted perturbations. Moreover, they propose some methods to mitigate this issue and enhance the model’s robustness.
Liu et al. (2023) extend the HumanEval dataset with more test cases using ChatGPT and mutation, and their findings revealed that a few test cases alone are inadequate to assess the correctness of a function. The inclusion of additional generated test cases led to a higher occurrence of failures, highlighting the importance of employing a diverse set of test cases to comprehensively evaluate function accuracy.
Besides, recent studies have explored various aspects of code generation. Such as, Li et al. (2024a) and Li et al. (2024b) have focused on secure code generation, while Ling et al. (2024) investigate social bias in code generation. In addition, Lin et al. (2024) examine code generation through the lens of emulating the software development process, and Saha et al. (2024) propose methods for generating code using the source code from other programming languages.
3 Methodology
We applied semantic-preserving perturbations to two types of coding tasks: (1) code generation given a Docstring and method signature, and (2) code completion given a Docstring and an incomplete method with its first few lines. For code generation task, we include the perturbations to revise the docstrings and method signatures. and in the code completion task, from docstrings and incomplete methods. We applied two types of function and Docstring perturbations for the code generation task, and format and syntax perturbations for the code completion task.
As our study focuses on comparisons across multiple programming languages, we prepared a robustness framework that can be applied to multiple PLs, including Java, C++, and JavaScript. In this section, we will discuss the methodology step by step.
3.1 Study Overview
This work focuses on the end-to-end code generation task under a black box setting, where it has access only to the model’s predictions with corresponding inputs. Figure LABEL:fig:overview illustrates the overall process of our study. We start with constructing the benchmark dataset for our experiments. Then, we apply 28 perturbation strategies from  Wang et al. (2022), focusing on four key areas in code. For code generation, we use perturbations on function names and docstrings, while for code completion, we apply them to format and syntax. These natural perturbations preserve the semantic meaning of the original prompts while reflecting common typographical errors developers may unintentionally introduce.
Finally, the original and the perturbed prompts are fed into the code generation model for evaluation separately. As we discussed earlier, the model takes a prompt as input and generates compilable code from the prompt. The prompt consists of a method signature, parameters, natural language comments, and others. In this work, we use an incomplete method as an input prompt, which consists of a docString and a method signature, which is identified as a nominal prompt throughout this work.
Note that this work does not cover adversarial attacks and only considers the robustness testing of the code generation models by generating perturbations randomly, providing them as input to the code generation models, and checking the performance with evaluation metrics.
The rest of this section discusses the dataset preparation, and the four types of perturbations, which consider DocString, function name, format, and syntax to generate examples, as well as the metrics used for robustness evaluation.
3.2 Dataset Preparation
In this part, we describe the steps taken to prepare the datasets used for evaluating our experiments. Our modifications aim to enhance the robustness evaluation of code generation models by extending the scope of assessment beyond previous benchmarks.
To address the limitations of Wang et al. (2022), which only evaluates code generation robustness in a single programming language (i.e., Python), we extend the evaluation to a multilingual setting by utilizing the HumanEval-X dataset. HumanEval-X, introduced by , serves as a benchmark for assessing the multilingual capabilities of code generation models.
Prior research (Liu et al., 2023) has highlighted critical shortcomings in the HumanEval dataset used by ReCode (Wang et al., 2022), particularly regarding the insufficient number of test cases per coding problem and imprecise natural language (NL) descriptions of expected program behaviors. Our analysis of HumanEval-X aligns with these findings, revealing that certain canonical solutions remain vulnerable to corner cases and that many test cases fail to adequately address such scenarios.
To improve evaluation robustness, we extend HumanEval-X by incorporating the enlarged test set provided by EvalPlus (Liu et al., 2023), which enhances HumanEval by refining NL descriptions, correcting solutions, and increasing the number of test cases per problem. Since EvalPlus is limited to Python, we converted its test cases into C++, Java, and JavaScript to facilitate multilingual evaluation. We also modified the canonical solutions in HumanEval-X to ensure compatibility with the additional test cases from EvalPlus, thereby enhancing the dataset’s comprehensiveness and robustness in evaluating code generation models.
We extend this dataset to multilingual by manually rewriting the prompts, canonical solutions, and test cases in C++, Java, and JavaScript through the following steps:
  1. (1)
    We find it necessary to update certain canonical solutions, as they are unable to handle corner cases effectively. To address this, we adopt the approach outlined in EvalPlus to implement the necessary improvements in the solutions. After making these updates, we proceed to create the prompts using the revised versions of the solutions.
  2. (2)
    We convert the test cases from EvalPlus, originally written in Python, into Java, C++, and JavaScript. During this process, we ensure that data types are transformed into their equivalent counterparts (e.g., converting lists to ArrayList in Java and lists to vector in C++).
  3. (3)
    We exclude test cases that are not compatible with Java and C++ (e.g., cases with mixed lists of integers, strings, and lists). Additionally, we attempt to modify the existing canonical solutions in Java, C++, and JavaScript to accommodate the converted test cases. However, if the updated solutions are unable to handle the test cases effectively, we exclude those specific test cases from the evaluation. For instance, Python allows for unlimited integer sizes, and some EvalPlus test cases tailored for Python might have very large integer values, surpassing the capacity of data types used in Java, C++, and JavaScript. Despite this challenge, we strive to maintain the method signature, including return types and parameter types, identical to those found in the original prompts.
3.3 Semantic Preserved Perturbation
After preparing our dataset by converting the test cases to multiple programming languages and updating the canonical solutions accordingly, we apply semantically preserved perturbations to the prompts for each language in a distinct manner. For the code generation task, we perturb the prompt by applying eight function name perturbations and ten docstring perturbations. For the code completion task, we first construct incomplete code snippets using the prompt and a portion of the canonical solution. We then apply seven syntax perturbations and six formatting perturbations to these incomplete codes. The specific details of the perturbation techniques are described in the following sections.
3.3.1 Perturbation on DocString
We applied 10 semantic-preserved perturbations on the DocString parts of the prompts. A docString is the human-written natural language to describe the method’s functionality. In a prompt, DocString provides the major instruction about how a method will be generated by a code generation model.
In this work, we follow ReCode’s implementation, which is based on NL-Augmenter (Dhole et al., 2021) for natural language transformations of datasets dealing with natural language [^2] https://github.com/GEM-benchmark/NL-Augmenter. We follow ten perturbation techniques for DocString from ReCode, which cover character-level, word-level, and sentence-level changes. As DocString is agnostic across languages, we directly apply the perturbations from ReCode to Java, C++, and JavaScript. The explanations with examples are shown in Table 1. These examples are collected from our perturbed dataset which we generated from the HumanEval-X dataset [^3] https://huggingface.co/datasets/THUDM/humaneval-x. For SynnonymSubstitution in DocString, synonyms by (Miller, 1995) are used to replace words in DocString with their synonyms. For BackTranslation, the whole DocString is translated into another language and then retranslated to English again to paraphrase (Sugiyama and Yoshinaga, 2019; Li and Specia, 2019).
Adaptation of Docstring Perturbations to multilingual To perturb a DocString, first, it is separated from the whole prompt, and the perturbation techniques are applied. The DocString may contain keywords of programming languages from code that must be untouched as changing them leads to altering the meaning. To maintain this, a blacklist of different programming languages is created which prevents the system from perturbing them. To separate the keywords and generate the blacklist, we utilize tree-sitter-java, tree-sitter-cpp, and tree-sitter-javascript for Java, C++, and JavaScript, respectively. The function names, variable names, and types are extracted from a code snippet by Tree-Sitter.
Table 1: Perturbations Examples for docString
Perturbation Explanation Example
Nominal Original Prompt Filter an input list of strings only for ones that start with a given prefix.
Back Translation Translate to another language and re-translate to English Filter an input list of strings only for strings starting with a certain prefix.
Butter Fingers Perturbation Replace random characters with some other characters Filher an input list of strings only for ones that start qith a given prefix.
Change Char Case Change case of random characters. FilTer aN inpuT list of strings Only fOr oneS that stArt With a GivEN prefix.
English Inflectional Variation Change the form of random verbs in a sentence Filter an input list of strings only for ones that start with a gives prefix.
Swap Characters Perturbation Swap two adjacent characters with each other Filtre an input list fo strings only for ones that start with a given prefix.
Synonym Insertion Insert a word’s synonym adjacent to its position Filter an input stimulation list of strings only for ones that start get down with a given impart prefix.
Synonym Substitution Replace a word with its synonym Filter an input list of strings only for ones that start with a give prefix.
Tense Transformation Future Transform the tense to future Filter an input list of strings only for ones that will start with a will give prefix.
Tense Transformation Past Transform the tense of the sentence to past Filter an input list of strings only for ones that started with a gave prefix.
Whitespace Perturbation Insert or remove whitespaces at random places Filter aninpu t list of str ings only for ones that startwith a given pr e fix.
3.3.2 Perturbation on Function Name
With this scope of perturbation, the function name is changed slightly, alike the DocString. The six perturbations used for Function renaming are shown in Table 2 with examples and short explanations.
Among the six perturbations, InflectionalVariation and SynonymSubstituition are word-level, and the rest of them are character-level perturbations. CamelCase perturbation is applied to the C++ dataset as the function names in HumanEval-X for C++ are written following the snake case format. On the other hand, SnakeCase is applied to Java and JavaScript function names as they follow the camel case format in function names.
Adaptation of Function Name Perturbations to Multiple Programming Languages: The semantic-preserved perturbations applied to function names require splitting a function name into multiple words. Different programming languages have different naming conventions. To apply semantics-preserving perturbations to function names, we first split the names into individual tokens based on the naming convention used in each language. For Java and JavaScript, we assume camelCase and split accordingly. For C++, we assume snake_case. If the detected naming style does not match the expected convention, it is first converted to the appropriate form (camelCase or snake_case) before splitting. Once split, the function name is treated as a sequence of words, and perturbations are applied to these words. These changes preserve the overall meaning while introducing variation in the name.
Table 2: Perturbations Examples for function name
perturbation explanation Example
Original original function name getPositive
SnakeCase change to snake cases from snake cases get_positive
CamelCase change to camel cases from camel cases getPositive
ButterFinger change a random character inside a function name getOositive
SwapChar swap two adjacent characters inside a function name getPsoitive
ChangeChar Change Case of random characters getposiTive
InflectionalVariation Change form of a word in a function name gottenPositive
SynonymSub Replace a word in a function name catchPositive
3.3.3 Perturbation on Syntax
Following the approach of Wang et al. (2022), we apply syntax-level perturbations to the code portion of prompts in a code completion setting. Unlike code generation, which uses only the docstring and method signature, code completion provides part of the function body to simulate a developer writing or editing code. Specifically, for each coding problem, we start with the canonical solution and select approximately the first half of its lines, following the strategy used in Wang et al. (2022), to create a baseline prompt. This prompt includes the original docstring and method signature, along with the selected code lines. We apply syntax-preserving perturbations only to these added lines. These perturbations introduce surface-level variation without changing the underlying semantics. To create partial code prompts for all three programming languages, we use the HumanEval-X dataset. For each language, we extract the canonical solution, truncate the function body, and apply the perturbation strategy described above. In this context, partial code refers to a prompt that includes the docstring, method signature, and a perturbed portion of the function body, tailored for code completion tasks.
We use six perturbations from ReCode (Wang et al., 2022) where three of them: DeadCode Insertion, Operand Swap, and For-While-Switch are generated by NatGen (Chakraborty et al., 2022). Additionally the rest three of them: Variable Renaming by CodeBERT (Feng et al., 2020), Variable Renaming with NatGen style, and generate random variable names using a combination of alphabetic and numeric characters are prepared by ReCode itself.
The following list explains the syntax perturbations used in this work. Table 4 illustrates examples of how the perturbations affect the original code snippet. The changes that happen in each of the syntax perturbations are highlighted.
  1. (1)
    DeadCodeInserter: Introduces irrelevant and non-executable code into the original codebase by inserting a code block at a random location. The inserted block may take the form of a loop that iterates zero times or an “if” statement that always evaluates to false. Its content is randomly selected from nearby code statements, constrained by small tree-sitter node sizes. This transformation is implemented using NatGen (Chakraborty et al., 2022), which provides mechanisms for programmatically modifying code syntax trees to inject such dead code in a syntactically valid way.
  2. (2)
    ForWhileTransformer: Select a random for loop and transfer it to an equivalent while loop or vice versa.
  3. (3)
    OperandSwap: Randomly pick a binary logical operation from a set of possible ones and switch the positions of the two operands involved in that operation (e.g. $a<b$ to $b>a$). If necessary, the operator is adjusted to ensure that the semantic meaning of the expression remains the same.
  4. (4)
    VarRenamerCB: This transformation substitutes a new name created by CodeBERT (Feng et al., 2020) for the most frequently used variable name in a particular code snippet. A mask token is used to substitute the original variable name. After that, CodeBERT generates a list of alternative names for each occurrence of the mask token along with a probability score. The name that received the highest rating overall gets chosen as the replacement.
  5. (5)
    VarRenamerNaive: This modification chooses the partial code’s most often used variable name and changes it to “VAR_0.” This is the NatGen package’s original implementation. The change does not alter the meaning.
  6. (6)
    VarRenamerRN: Selects the most commonly used variable name from the partial code and replaces it with a randomly generated string composed of a mix of alphabetic and numeric characters.
Adaptation of Syntax Perturbations to multilingual: Initially, the task involves generating partial prompts from the nominal prompts. To accomplish this, we make adjustments to ReCode’s code and introduce specific rules for creating partial code, effectively separating the header, DocString, body, and other segments for Java, C++, and JavaScript. Then, we leverage the NatGen libraries to generate syntax-level perturbations.
Figure 4: Syntax Perturbations (JavaScript)
3.3.4 Perturbation on Format
This perturbation is employed to evaluate whether the code generation model can effectively handle a perturbed prompt with identical information to the original one but with a distinct formatting style. It tests the model’s ability to handle variations in formatting while generating code. In this work, we use 6 types of formats as follows.
  1. (1)
    doc2comments: This perturbation transforms a Docstring (written as a block comment) into multiple lines of inline comments.
  2. (2)
    new_line_aftercode: A new empty line is added after the code.
  3. (3)
    new_line_afterdoc: A new empty line is added after the docString.
  4. (4)
    new_lines: A new line is added at a random position between any two lines in the full prompt.
  5. (5)
    split_lines: Split the longest line of code and split it into two lines in the middle.
  6. (6)
    tab_indent: random replacements of space indentation with tabs or replace tabs with four spaces. Although languages such as Java, JavaScript, and C++ are not sensitive to indentation, this perturbation is used to check the robustness of code generation models.
[…]
By now, we have proposed a total of 28 strategies for perturbation. And in our experiments, for each coding problem, we apply each of the 28 strategies to the original prompt and generate five perturbed prompts. Therefore, for each coding problem, in total, we have 28*5 perturbed prompts, and for each perturbed prompt, one code is generated.
3.4 Evaluation
There are several metrics used by previous works to evaluate the performance of code generation. In our evaluation, we consider both the worst-case passes (i.e., for each coding problem, all perturbed prompts $s$ of a prompt must be passed) and partial-case passes (a partial number $t<s$ of the perturbed prompts from a prompt must be passed). Alike ReCode (Wang et al., 2022), we use the following metrics for our evaluation. We choose a greedy approach for all of the metrics, skipping the sampling process to eliminate the randomness effect and enable fair comparisons (Wang et al., 2022).
3.4.1 Robust Pass ($RP_{s}@k$)
Pass@k is a commonly used metric for code generation tasks. As we are generating multiple perturbed prompts from a single prompt, we use a modified version of this metric. From the initial prompt, there is $s$ number of perturbed prompts. Additionally, $k$ generations are made from each perturbed prompt, for a total of $s*k$ generations from a single prompt. In this work, we use $s=5$ and $k=1$. We use the $RP_{s}@k$ metrics as equation 2.
$RP_{s}@k=\frac{RC_{s}}{N}$ (2)
where $RC_{s}$ denotes the number of correct prompts and $N$ denotes the total number of prompts.
3.4.2 Robust Drop ($RD_{s}@k)$
To compare the robustness of the generated code from perturbed data with the generated code from nominal data, the Robust Drop metric is used. This metric states the fall in the performance compared to the nominal data. In equation 3, this metric is defined.
$RD_{s}@k=\frac{Pass_{nominal}@k-RP_{s}@k}{Pass_{nominal}@k}$ (3)
where $Pass_{nominal}@k$ denotes the result for the nominal dataset and $RP_{s}@k$ denotes the robust relative value for the perturbed dataset we get from equation 2.
3.4.3 Robust Relative ($RR_{s}@k)$
The Robust Drop metric provides an idea of the performance decrease for the perturbed dataset. But it does not highlight the pair-by-pair performance change between nominal and perturbed datasets. To get the relative changes, the Robust Relative metric is used. There can be instances where models produce inaccurate code for the nominal prompts but accurately predict the perturbed ones. This non-robust behavior should also be taken into account when reporting the model’s robustness. For this situation, the Robust Relative metric is useful. In equation 4, the Robust Relative metric is defined.
$RR_{s}@k=\frac{RC_{s}^{[+]}+RC_{s}^{[-]}}{N}$ (4)
where $RC_{s}^{[+]}$ denotes the number of correct cases in perturbations that were detected as incorrect cases for the original prompts. $RC_{s}^{[-]}$ denotes the opposite, the number of incorrect cases in perturbations that were detected as correct cases for the original prompts. This work follows a greedy setting (n=1 sample), and the equation is made upon this setting, and this metric denotes the number of positive or negative changes for a set of prompts.
4 Experiment Setup
In this section, we present the experimental design and setup employed to evaluate the robustness of code generation models. By meticulously addressing each sub-section, we provide a comprehensive overview of our experiments, enabling a deeper understanding of our research implementation and subsequent analysis.
4.1 Dataset
As stated in Section 3.2, we utilized two datasets to evaluate the robustness of code generation models. Below, we further detail the datasets used in our experiments.
HumanEval-X. HumanEval-X [^4] https://huggingface.co/datasets/THUDM/humaneval-x is an extended version of the HumanEval [^5] https://github.com/openai/human-eval dataset, originally designed to evaluate code generation models in Python. HumanEval-X expands this benchmark to support multiple programming languages, including C++, Java, JavaScript, and Go. It consists of 164 prompts, each accompanied by a canonical solution and language-specific test cases. In our study, we leverage HumanEval-X to assess code translation and bug classification across Java, C++, and JavaScript. The dataset provides a total of X test cases, with an average of Y test cases per language, ensuring comprehensive evaluation across different programming paradigms.
EvalPlus. As the second dataset, we use EvalPlus after the curation process described in Section 3.2. The prompts remain consistent with those in HumanEval-X, except for cases where we made adjustments to accommodate corner cases. Additionally, EvalPlus provides more extensive and critical test cases. However, many HumanEval-X solutions were not directly suitable for EvalPlus test cases. To address this, we manually wrote canonical solutions for each programming language in HumanEval-X and then tested them using the EvalPlus-converted test cases. This adaptation ensures that the evaluations are meaningful across multiple programming languages. The statistics of the test cases are shown in Table 3
Table 3: Statistical measures of the number of test cases for different languages
Avg Median Std
Language HumanEval-X Ours HumanEval-X Ours HumanEval-X Ours
CPP 6.96 722.56 7.0 973.5 3.73 375.95
Java 6.5 748.48 6.0 1000.0 4.55 359.39
JS 7.95 709.99 8.0 972.0 3.73 379.72
4.2 Studied Code Language Model
In our experiment, we focus on evaluating four distinct code generation models. Specifically, we select Incoder-1b, Incoder-6b, CodeGen-2b-Multi, and CodeGen-6b-Multi as they have been trained in multiple programming languages. These models are specifically chosen due to their varying number of parameters, allowing us to assess our experiments across a range of model complexities. Due to the excessive size and our inability to run them on our machine, we exclude certain large models, such as CodeGen-16B-Multi, from our experiment.
4.3 Setup
[…]
4.4 Evaluation of Perturbation Quality
As stated in Section 3, the perturbations should preserve the semantic meaning of the original prompts while appearing natural. To evaluate the perturbed dataset’s quality, we employ three measurements for the docString, function name, and code sections described in the following.
Table 4: Similarity Values Between Nominal and Perturbed Datasets
Scope Java CPP JavaScript ReCode (Python)
Syntax 0.96 0.97 0.95 0.96
DataFlow 0.97 0.97 0.94 0.97
DocString 0.93 0.93 0.93 0.93
Func Name 0.79 0.85 0.79 0.81
4.4.1 Sentence Similarity Measurement
For the Java, C++, and JavaScript datasets, we assess the semantic similarity between the perturbed and non-perturbed prompts. To do this, we employ the sentence cosine similarity measurement, following the same methodology as ReCode. We calculate the similarities between the perturbed and nominal datasets using the sentence embeddings from MPNet (Song et al., 2020). Specifically, we utilize MPNet embeddings to measure the cosine similarities for the docString and function name sections. For docstrings, we directly compare them after extraction, while for function names, we split them into lists of words and then compare them. Table 4 shows the summary of similarities, and Table 5 shows the similarities for individual perturbations.
Table 5: Sentence Similarities for individual docString and function name perturbations
Categories Perturbations Java JavaScript CPP docString ChangeCharCase 1.0 1.0 1.0 TenseTransformationPast 0.98 0.98 0.98 EnglishInflectionalVariation 0.95 0.95 0.95 WhitespacePerturbation 0.89 0.9 0.89 ButterFingersPerturbation 0.86 0.86 0.87 TenseTransformationFuture 0.97 0.97 0.97 SynonymSubstitution 0.88 0.88 0.88 SwapCharactersPerturbation 0.9 0.89 0.9 SynonymInsertion 0.9 0.89 0.91 BackTranslation 0.94 0.93 0.92 Function FuncRenameSynonymSub 0.72 0.75 0.72 FuncRenameButterFinger 0.56 0.62 0.57 FuncRenameSwapChar 0.83 0.83 0.82 FuncRenameSnake/CamelCase 0.89 0.89 0.89 FuncRenameInflectionalVariation 0.93 0.94 0.93 FuncRenameChangeChar 0.37 0.41 0.36
4.4.2 Code Similarity Measurement
To assess the similarities between the perturbed and nominal data, we utilize CodeBLEU’s Syntax and DataFlow scores (Ren et al., 2020). For our analysis, we adapt the CodeBLEU score calculation method employed by CodeXGlue (Lu et al., 2021). Additionally, to calculate scores for C++, we incorporate NaturalCC (Wan et al., 2022) since the DataFlowGraph for C++ is not available in CodeXGlue, which is essential for computing the DataFlow score. The summary of similarity scores is presented in Table 4, while Table 6 displays the individual similarity scores. In general, the scores are higher, indicating significant similarity, with the exception of DeadCodeInserter. The inclusion of a code chunk within the code block leads to the dissimilarity between the perturbed prompt and the nominal prompt.
Table 6: CodeBLEU Similarities for individual perturbations
Categories Perturbations Java CPP JavaScript Syntax Dataflow Syntax Dataflow Syntax Dataflow Syntax ForWhileTransformer 0.94 0.99 0.96 1.0 0.93 0.99 OperandSwap 0.87 0.99 0.97 1.0 0.87 1.0 DeadCodeInserter 0.76 0.75 0.76 0.64 0.74 0.72 VarRenamerRN 0.99 0.98 0.99 1.0 1.0 0.98 VarRenamerCB 0.98 0.97 1.0 1.0 0.99 0.97 VarRenamerNaive 0.99 0.99 0.99 1.0 1.0 0.98 Format newLineAfterdoc 1.0 1.0 1.0 1.0 1.0 1.0 SplitLines 0.99 1.0 0.99 1.0 0.97 1.0 Doc2Comments 1.0 1.0 0.92 0.99 0.89 0.69 NewLineAftercode 1.0 1.0 1.0 1.0 1.0 1.0 NewLines 1.0 1.0 1.0 1.0 1.0 1.0 TabIndent 1.0 1.0 1.0 1.0 1.0 1.0
4.4.3 Human Annotation of Naturalness and Semantic Similarity in Perturbed Prompts
We conduct a random sampling of 300 perturbed and non-perturbed pairs, comprising 100 Java, 100 C++, and 100 JavaScript prompts. Each prompt is annotated by a single annotator, who provides three values: the naturalness of the perturbed prompt, the naturalness of the non-perturbed prompt, and the semantic similarity between the perturbed and non-perturbed prompts. The naturalness score ranges continuously from 0 to 1, where 0 indicates unnatural text and 1 signifies highly natural text. Similarly, the semantic similarity score also ranges from 0 to 1, where 0 represents a complete change in meaning, 0.5 indicates partial changes, and 1 denotes exact semantic preservation. We adopted the annotation methodology employed by ReCode (Wang et al., 2022), which utilized five annotators. However, for our study, we used one annotator to annotate the 300 samples. The findings of the annotations are presented in Table 7. Table 7 summarizes the human annotation results across Java, C++, and JavaScript, and also includes a comparison with ReCode’s Python-based annotations. The results show that non-perturbed prompts maintain near-perfect naturalness scores (above 0.98), while perturbed prompts still retain relatively high naturalness (around 0.8–0.85). Semantic similarity scores are also consistently high (above 0.94), indicating that the perturbations generally preserve meaning. These findings suggest that although the perturbations slightly reduce naturalness, they largely maintain the original semantics, validating the use of these perturbations in robustness evaluations.
Table 7: Result of Human Annotations for 300 Random Samples in Three Different Languages
Scope Java CPP JavaScript ReCode (Python)
Naturalness (Nominal) $\uparrow$ 0.99 0.98 0.98 0.92
Naturalness (Perturbed) $\uparrow$ 0.85 0.81 0.8 0.75
Semantic Similarity $\uparrow$ 0.94 0.96 0.95 0.92
5 Experiment Results
In this section, we aim to answer three research questions (RQs). For each RQ, we present the motivation and the approach to addressing the RQ, and the corresponding results.
RQ1: How robust are large language models (LLMs) across different programming languages?
Motivation: Our experiment encompasses three distinct languages and four different models. A prior investigation (Wang et al., 2022) demonstrates the influence of perturbations on the robustness of code generation models. Nonetheless, the question of whether language-specific variations impact robustness remains unanswered. Hence, in this study, we aim to address this gap by exploring discrepancies in Large Language Models’ (LLMs) robustness across diverse programming languages (Java, C++, and JavaScript), shedding light on the extent of disparities among these languages.
Approach: To address this research question, we perform experiments using Java, C++, and JavaScript across four models: Incoder-1B, Incoder-6B, CodeGen-Multi-2B, and CodeGen-Multi-6B. We evaluate the performance using three metrics: $RP_{s}@k$ measures robustness, $RD_{s}@k$ assesses the decrease in robustness due to changes, and $RR_{s}@k$ captures the relative differences between the original and modified datasets. The perturbations related to the Function Name and the DocString are compared with the nominal prompts (prompts that have only DocString and method signature), while the perturbations related to Syntax and Format are compared with the partial prompts (prompts that have some lines of code along with DocString and method signature)
Results: Overall, the performance of models deteriorates after applying the perturbations. Table 8 presents the worst-case performance of four models—Incoder-1B, Incoder-6B, CodeGen-2B-Multi, and CodeGen-6B-Multi—under different perturbation types (DocString, Function Signature, Syntax, and Format) across three programming languages (Java, C++, and JavaScript). Similarly, Table 9 shows the results in partial-case scenarios, where a model passes at least 3 out of 5 generated samples.
Across both tables, we observe a consistent performance drop after applying perturbations. For instance, in Table 8, the metric $\text{RD}_{5}@1$, which quantifies the worst-case degradation, shows severe drops—up to 0.9 (90%) in some DocString perturbation cases for Incoder-1B (Java). Even more superficial perturbations, such as Syntax and Format, lead to noticeable performance degradation, albeit to a lesser extent. Table 9 highlights that even under relaxed partial-pass criteria, the relative performance drop (RR5@1) remains substantial, typically ranging from 5% to 20%, depending on the model and language. This suggests that perturbations—whether semantic or syntactic—can significantly impact model behavior and reliability.
Interestingly, we observe a slight performance improvement in the partial-case scenario for Java under DocString and Function perturbations. This suggests that in some cases, modifying the function or the NL description can yield better results, hinting at the potential of prompt refinement as a strategy for improving code generation performance.
In general, all models demonstrate performance degradation under perturbations, reaffirming the brittleness of current LLM-based code generation systems. While models such as CodeGen-6B show slightly better robustness, the consistent drop across all settings underscores the need for robustness-aware evaluation and training. Moreover, the variation in impact across programming languages suggests that the models’ sensitivity may be language-dependent, warranting further investigation into language-specific robustness strategies.
If we see the language-wise trends, the three languages—Java, C++, and JavaScript—exhibit distinct behaviors under input perturbations. Java shows the highest robustness, with notable RP5@1 gains and minimal RD5@1 drops, particularly under Format and Syntax perturbations. Its structured syntax likely helps models infer intent despite surface-level changes.
C++ is the most fragile across the board. It suffers large RD5@1 drops, especially under worst-case perturbations, with only modest improvements under partial conditions. Its complex syntax and variability appear to challenge the models more than the other languages.
JavaScript falls in between. It benefits from Format and Syntax perturbations with decent RP5@1 gains, though it still shows some drops in DocString and Function changes. Its flexible syntax likely helps models recover better than in C++ but not as effectively as in Java.
Overall, Java is the most resilient, followed by JavaScript, while C++ is the most sensitive to input changes.
A Larger model size does not always guarantee better robustness or pass rate across all programming languages. As shown in Table 8 and Table 9, in general, we can observe a robustness or pass rate improvement with the increase of model sizes. For instance, for CodeGen in the worst case scenario, the $RD_{5}@1$ of JavaScript increases from 0.07 to 0.08 for 2B and 6B, respectively. On the contrary, for Java, we also observe that there is a drop of 0.02 of the $RD_{5}@1$. The results show that increasing the model size does not always guarantee better robustness and highlight the need for more comprehensive experiments on different programming languages.
CodeGen has a relatively higher pass@k than Incoder across different languages. By checking the Nominal metric in Table 8 and Table 9, with the same parameter size (e.g., 6B), CodeGen usually has a better performance than that of Incoder. For example, for Java, CodeGen has a pass@k of 0.16 compared to 0.11 produced by Incoder. This might be due to the different sizes of tokens used for the training, as CodeGen has been trained on more than 1000 billion tokens and Incoder has only been trained on about 60 billion tokens. The other reason may come from the distribution of the training data. Although the training data of Incoder contains 28 languages, which is much more diverse than that of CodeGen (i.e., trained in six programming languages), the majority of the data is Python and JavaScript. This may also explain why Incoder only has the worst pass@k of 0.02 on the Java coding problems.
Table 8: Performance summary of perturbations against code generation models using HumanEval-X test cases and worst case passes (5 out of 5). Up arrows indicate the extent of performance improvement, while down arrows indicate the extent of performance drop.
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS
[…]
RP5@1$\uparrow$ 0.0 0.04 0.03 0.04 0.06 0.05 0.07 0.07 0.05 0.05 0.08 0.06
RD5@1$\downarrow$ 0.9 0.57 0.63 0.68 0.49 0.58 0.45 0.43 0.59 0.68 0.44 0.65
RR5@1$\downarrow$ 0.02 0.08 0.05 0.09 0.06 0.08 0.06 0.06 0.08 0.11 0.07 0.11
[…]
RP5@1$\uparrow$ 0.0 0.04 0.04 0.04 0.07 0.07 0.08 0.07 0.07 0.05 0.08 0.07
RD5@1$\downarrow$ 0.88 0.57 0.49 0.64 0.46 0.47 0.4 0.44 0.48 0.69 0.4 0.56
RR5@1$\downarrow$ 0.02 0.07 0.05 0.08 0.06 0.07 0.05 0.06 0.07 0.11 0.06 0.11
Syntax Nominal (Partial)$\uparrow$ 0.33 0.1 0.38 0.45 0.17 0.44 0.56 0.13 0.39 0.54 0.22 0.44
RP5@1$\uparrow$ 0.23 0.03 0.23 0.27 0.05 0.27 0.33 0.07 0.2 0.29 0.09 0.23
RD5@1$\downarrow$ 0.3 0.73 0.39 0.38 0.68 0.39 0.41 0.45 0.49 0.46 0.58 0.47
RR5@1$\downarrow$ 0.21 0.08 0.16 0.24 0.14 0.17 0.26 0.08 0.2 0.25 0.14 0.24
Format Nominal (Partial)$\uparrow$ 0.33 0.1 0.38 0.45 0.17 0.44 0.56 0.13 0.39 0.54 0.22 0.44
RP5@1$\uparrow$ 0.2 0.04 0.26 0.28 0.07 0.3 0.45 0.1 0.27 0.42 0.12 0.27
RD5@1$\downarrow$ 0.38 0.64 0.33 0.36 0.57 0.32 0.2 0.22 0.31 0.21 0.44 0.39
RR5@1$\downarrow$ 0.16 0.07 0.14 0.19 0.11 0.15 0.13 0.04 0.16 0.13 0.1 0.21
Table 9: Performance summary of perturbations using partial cases using HuamEval-X test cases and partial passes (3 out of 5)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS
[…]
RP5@1$\uparrow$ 0.03 0.07 0.07 0.07 0.11 0.11 0.12 0.12 0.1 0.12 0.13 0.12
RD5@1$\downarrow$ -0.07 0.31 0.16 0.35 0.11 0.16 0.09 0.05 0.23 0.22 0.07 0.28
RR5@1$\downarrow$ 0.02 0.06 0.05 0.06 0.04 0.05 0.04 0.05 0.05 0.08 0.04 0.09
[…]
RP5@1$\uparrow$ 0.03 0.08 0.08 0.07 0.1 0.12 0.12 0.12 0.11 0.11 0.12 0.15
RD5@1$\downarrow$ -0.04 0.19 -0.01 0.37 0.15 0.1 0.06 0.05 0.11 0.28 0.12 0.12
RR5@1$\downarrow$ 0.02 0.05 0.04 0.07 0.05 0.05 0.05 0.05 0.05 0.09 0.04 0.07
Syntax Nominal (Partial)$\uparrow$ 0.33 0.1 0.38 0.45 0.17 0.44 0.56 0.13 0.39 0.54 0.22 0.44
RP5@1$\uparrow$ 0.27 0.07 0.28 0.36 0.11 0.32 0.4 0.11 0.28 0.39 0.14 0.33
RD5@1$\downarrow$ 0.18 0.37 0.26 0.19 0.36 0.27 0.29 0.13 0.28 0.27 0.35 0.24
RR5@1$\downarrow$ 0.2 0.08 0.11 0.2 0.17 0.13 0.21 0.09 0.16 0.18 0.13 0.17
Format Nominal (Partial)$\uparrow$ 0.33 0.1 0.38 0.45 0.17 0.44 0.56 0.13 0.39 0.54 0.22 0.44
RP5@1$\uparrow$ 0.32 0.07 0.32 0.39 0.13 0.41 0.52 0.14 0.36 0.53 0.19 0.35
RD5@1$\downarrow$ 0.03 0.29 0.18 0.12 0.21 0.06 0.07 -0.07 0.09 0.01 0.12 0.2
RR5@1$\downarrow$ 0.11 0.05 0.09 0.14 0.1 0.07 0.08 0.04 0.12 0.05 0.06 0.15
RQ2: What are the major factors that contribute to the difference in the robustness of LLMs on different programming languages?
Motivation: From the insights of RQ1, it becomes apparent that the performance of identical code generation models varies across different languages. This research inquiry delves deeper into uncovering the underlying reasons for such disparities. By doing so, we aim to find out the pivotal factors that contribute significantly to the fluctuations in robustness observed among various programming languages. Through this exploration, we seek to shed light on the influential components that drive the differences in how well models perform on perturbed data across different linguistic contexts.
Approach: To answer this research question, we first find out the error rates for the nominal, partial, and each category of perturbations. We categorize the run statuses as passed, assertion error, compilation error, timeout, and runtime error. To calculate these ratios, we quantify the total occurrence of each category and divide it by the overall number of instances, thereby providing insights into the distribution of these outcomes. Subsequently, we engage in a manual analysis of select samples from each programming language, delving into the underlying causes of failures and variations in robustness across different languages.
Results: The resultant error rates are presented in Table 10. From the table, it is observed that nominal prompts exhibit similar pass rates across the three languages. Furthermore, the pass rates for perturbations stemming from nominal prompts display analogous patterns. On the other hand, notable differences emerge in the error rates associated with partial data. In partial cases, C++ shows a lower pass rate compared to Java and JavaScript. Notably, C++ yields a significantly higher count of assertion errors in comparison to the other two languages, suggesting that C++ fares less favorably in code completion tasks when contrasted with Java and JavaScript.
Table 10: Error Percentages for code generation models
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Type Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal Passed 2.44 10.37 7.93 10.98 12.2 12.8 12.8 12.2 12.8 15.85 14.02 16.46
Assertion 31.1 60.37 55.49 55.49 62.8 63.41 69.51 68.29 60.98 66.46 67.07 48.78
Compilation 58.54 29.27 35.98 26.83 22.56 23.17 9.15 18.29 25.61 5.49 17.07 33.54
Timeout 1.83 0.0 0.61 0.61 0.61 0.61 0.61 0.61 0.61 1.22 1.22 1.22
Runtime 6.1 0.0 0.0 6.1 1.83 0.0 7.93 0.61 0.0 10.98 0.61 0.0
DocString Passed 2.9 8.57 6.8 8.17 11.37 11.27 12.43 12.87 10.51 13.0 13.9 12.88
Assertion 33.3 55.95 49.07 51.13 60.74 62.72 65.2 63.46 49.71 57.46 65.6 42.85
Compilation 57.32 34.54 42.21 32.07 26.67 25.49 14.16 22.45 39.04 21.95 18.7 43.57
Timeout 0.95 0.67 1.91 1.04 0.44 0.52 1.0 0.78 0.74 1.37 1.11 0.7
Runtime 5.52 0.27 0.0 7.59 0.78 0.0 7.22 0.44 0.0 6.22 0.7 0.0
Function Passed 2.85 9.13 7.89 8.6 11.43 12.2 12.56 12.63 12.01 12.46 13.34 14.74
Assertion 34.47 59.09 54.55 55.06 64.39 64.43 68.43 67.3 56.1 61.4 69.58 47.36
Compilation 55.63 30.71 35.71 27.05 22.56 22.97 10.93 18.87 31.04 18.9 14.77 37.05
Timeout 1.06 0.89 1.85 0.83 0.52 0.41 1.02 0.71 0.85 0.77 1.36 0.85
Runtime 6.0 0.17 0.0 8.46 1.1 0.0 7.05 0.49 0.0 6.46 0.94 0.0
Partial Passed 32.93 10.37 38.41 44.51 17.07 43.9 56.1 12.8 39.02 53.66 21.95 43.9
Assertion 31.71 62.2 48.78 42.07 63.41 44.51 35.98 67.07 43.9 38.41 65.24 37.8
Compilation 30.49 26.22 12.8 9.15 17.68 11.59 5.49 17.07 17.07 4.27 10.98 17.68
Timeout 1.83 0.61 0.0 1.83 0.61 0.0 0.61 1.83 0.0 0.0 0.61 0.61
Runtime 3.05 0.61 0.0 2.44 1.22 0.0 1.83 1.22 0.0 3.66 1.22 0.0
Syntax Passed 27.66 7.95 28.64 36.67 12.32 32.78 40.59 12.09 29.07 39.09 14.8 33.46
Assertion 25.79 45.79 43.15 28.39 50.63 41.32 32.93 56.38 40.08 28.8 55.53 39.04
Compilation 41.97 44.92 27.28 31.4 35.69 24.84 22.28 30.18 29.94 28.05 28.17 26.99
Timeout 1.73 1.1 0.93 0.59 1.14 1.06 0.57 1.08 0.91 0.71 1.28 0.51
Runtime 2.85 0.24 0.0 2.95 0.22 0.0 3.64 0.26 0.0 3.35 0.22 0.0
Format Passed 31.04 8.68 31.75 39.19 14.76 38.78 52.36 14.61 35.59 52.42 19.37 35.63
Assertion 31.87 58.68 46.52 35.87 62.54 40.24 36.38 66.18 44.13 35.14 65.71 42.8
Compilation 33.15 31.18 21.02 20.55 20.51 20.61 7.72 16.95 19.72 9.19 12.66 21.36
Timeout 0.96 1.14 0.71 0.96 1.24 0.37 0.43 1.06 0.57 0.28 0.96 0.2
Runtime 2.99 0.33 0.0 3.43 0.96 0.0 3.11 1.2 0.0 2.97 1.3 0.0
RQ3: How the enriched test cases affect the performance of the LLMs under testing?
Motivation: The findings from a recent study (Liu et al., 2023) highlight the limitation of humaneval test cases which can not detect all of the incorrect samples. To enhance the precision and depth of our investigation, we use the translated test cases from evalplus for the evaluation of generated code in our study. By doing this, we aim to ensure a more thorough exploration of the errors generated by the models, enabling us to better uncover the scope of inaccuracies in their code generation among the three languages we consider.
Approach: To answer this research question, we follow the same procedure as we follow for HumanEval-X test cases. We use the same four models (Incoder1B, Incoder6B, CodeGen-Multi-2B, and CodeGen-Multi-6B) and three languages (Java, C++, and JavaScript), and the same metrics ($RP_{s}@k$, $RD_{s}@k$, and $RR_{s}@k$) to evaluate the performance.
Figure 5: A generated code by CodeGen-6B-Multi model which passes the test cases of HumanEval-X but fails against the evalplus translated test case
Results: Using more test cases results in reduced model robustness compared to using only the original test cases. Table 11 and Table 12 present the results for worst-case and partial-case scenarios using EvalPlus test cases. Analyzing these results alongside Table 8 and Table 9 reveals a consistent performance decline across all cells. This suggests that EvalPlus test cases cause more failures in samples that previously passed the HumanEval-X test cases. For instance, Figure 5 shows a code example generated from a Back-Translation perturbed prompt that successfully passes a HumanEval-X test case but times out on a new EvalPlus test case.
Table 11: Performance summary of perturbations against code generation models using EvalPlus test cases and worst case passes (5 out of 5)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS
[…]
RP5@1$\uparrow$ 0.0 0.03 0.02 0.03 0.05 0.04 0.06 0.05 0.03 0.05 0.05 0.05
RD5@1$\downarrow$ 0.87 0.71 0.58 0.74 0.51 0.57 0.51 0.6 0.65 0.65 0.56 0.61
RR5@1$\downarrow$ 0.02 0.07 0.04 0.08 0.05 0.06 0.07 0.07 0.06 0.09 0.07 0.07
[…]
RP5@1$\uparrow$ 0.0 0.02 0.03 0.03 0.05 0.04 0.07 0.05 0.04 0.04 0.07 0.05
RD5@1$\downarrow$ 0.89 0.71 0.46 0.71 0.47 0.59 0.42 0.56 0.61 0.68 0.4 0.53
RR5@1$\downarrow$ 0.02 0.07 0.04 0.08 0.05 0.06 0.05 0.07 0.06 0.1 0.06 0.07
Syntax Nominal (Partial)$\uparrow$ 0.32 0.08 0.35 0.43 0.13 0.41 0.55 0.12 0.36 0.54 0.2 0.41
RP5@1$\uparrow$ 0.22 0.02 0.22 0.26 0.04 0.25 0.32 0.05 0.19 0.28 0.07 0.22
RD5@1$\downarrow$ 0.3 0.81 0.39 0.39 0.7 0.4 0.42 0.62 0.46 0.48 0.63 0.47
RR5@1$\downarrow$ 0.21 0.07 0.14 0.24 0.11 0.16 0.25 0.09 0.17 0.26 0.14 0.23
Format Nominal (Partial)$\uparrow$ 0.32 0.08 0.35 0.43 0.13 0.41 0.55 0.12 0.36 0.54 0.2 0.41
RP5@1$\uparrow$ 0.2 0.03 0.24 0.28 0.06 0.29 0.44 0.08 0.26 0.41 0.1 0.26
RD5@1$\downarrow$ 0.38 0.65 0.33 0.35 0.58 0.3 0.2 0.38 0.27 0.23 0.51 0.38
RR5@1$\downarrow$ 0.16 0.06 0.13 0.18 0.08 0.14 0.13 0.05 0.14 0.13 0.1 0.2
Table 12: Performance summary of perturbations using partial cases using EvalPlus test cases and partial passes (3 out of 5)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS
[…]
RP5@1$\uparrow$ 0.02 0.05 0.05 0.06 0.08 0.08 0.11 0.09 0.07 0.11 0.1 0.08
RD5@1$\downarrow$ -0.07 0.46 0.02 0.38 0.07 0.17 0.12 0.18 0.27 0.19 0.19 0.27
RR5@1$\downarrow$ 0.02 0.06 0.04 0.06 0.03 0.04 0.04 0.05 0.05 0.05 0.04 0.06
[…]
RP5@1$\uparrow$ 0.02 0.06 0.06 0.06 0.08 0.08 0.11 0.1 0.08 0.1 0.11 0.1
RD5@1$\downarrow$ -0.11 0.27 -0.02 0.35 0.12 0.16 0.08 0.16 0.09 0.26 0.09 0.12
RR5@1$\downarrow$ 0.02 0.06 0.04 0.05 0.04 0.04 0.04 0.06 0.05 0.06 0.04 0.05
Syntax Nominal (Partial)$\uparrow$ 0.32 0.08 0.35 0.43 0.13 0.41 0.55 0.12 0.36 0.54 0.2 0.41
RP5@1$\uparrow$ 0.26 0.05 0.26 0.35 0.09 0.3 0.38 0.08 0.27 0.38 0.12 0.31
RD5@1$\downarrow$ 0.17 0.42 0.25 0.2 0.36 0.28 0.31 0.34 0.26 0.29 0.39 0.25
RR5@1$\downarrow$ 0.21 0.08 0.1 0.2 0.13 0.13 0.21 0.09 0.14 0.18 0.12 0.17
Format Nominal (Partial)$\uparrow$ 0.32 0.08 0.35 0.43 0.13 0.41 0.55 0.12 0.36 0.54 0.2 0.41
RP5@1$\uparrow$ 0.31 0.05 0.29 0.38 0.11 0.4 0.51 0.11 0.34 0.53 0.16 0.33
RD5@1$\downarrow$ 0.01 0.33 0.18 0.12 0.2 0.04 0.08 0.09 0.06 0.02 0.16 0.2
RR5@1$\downarrow$ 0.11 0.04 0.09 0.14 0.07 0.06 0.09 0.05 0.1 0.05 0.05 0.14
Table 13: Performance of code generation models against docstring perturbations (HumanEval-X and worst case passes)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS Nominal RP5@1 0.02 0.1 0.09 0.11 0.12 0.13 0.13 0.13 0.13 0.16 0.15 0.16 EnglishInflectionalVariation RP5@1 0.0 0.04 0.03 0.04 0.05 0.06 0.07 0.06 0.05 0.07 0.08 0.05 RD5@1 1.0 0.59 0.62 0.67 0.55 0.52 0.48 0.5 0.62 0.58 0.43 0.7 RR5@1 0.02 0.07 0.05 0.07 0.07 0.07 0.06 0.06 0.08 0.1 0.06 0.12 TenseTransformationFuture RP5@1 0.01 0.05 0.04 0.04 0.07 0.07 0.09 0.08 0.07 0.05 0.08 0.06 RD5@1 0.75 0.47 0.46 0.61 0.45 0.48 0.33 0.35 0.43 0.65 0.43 0.63 RR5@1 0.02 0.07 0.04 0.08 0.05 0.07 0.05 0.05 0.05 0.1 0.06 0.1 ButterFingersPerturbation RP5@1 0.0 0.04 0.04 0.03 0.06 0.05 0.07 0.06 0.04 0.04 0.07 0.06 RD5@1 1.0 0.59 0.46 0.72 0.5 0.57 0.48 0.5 0.67 0.73 0.48 0.63 RR5@1 0.02 0.09 0.05 0.09 0.06 0.07 0.06 0.07 0.09 0.12 0.07 0.1 ChangeCharCase RP5@1 0.0 0.01 0.01 0.02 0.06 0.04 0.05 0.07 0.03 0.04 0.07 0.04 RD5@1 1.0 0.88 0.85 0.78 0.5 0.71 0.62 0.45 0.76 0.73 0.48 0.74 RR5@1 0.02 0.09 0.08 0.1 0.07 0.09 0.08 0.05 0.1 0.12 0.07 0.12 BackTranslation RP5@1 0.0 0.04 0.04 0.04 0.08 0.07 0.09 0.09 0.08 0.06 0.09 0.07 RD5@1 1.0 0.59 0.54 0.61 0.35 0.43 0.33 0.3 0.38 0.62 0.39 0.56 RR5@1 0.02 0.09 0.04 0.08 0.05 0.08 0.07 0.04 0.06 0.1 0.07 0.1 SwapCharactersPerturbation RP5@1 0.0 0.04 0.02 0.02 0.07 0.05 0.09 0.07 0.05 0.04 0.07 0.07 RD5@1 1.0 0.59 0.77 0.78 0.45 0.57 0.33 0.45 0.62 0.73 0.48 0.59 RR5@1 0.02 0.09 0.06 0.09 0.05 0.09 0.05 0.05 0.08 0.12 0.07 0.11 TenseTransformationPast RP5@1 0.0 0.06 0.04 0.04 0.06 0.06 0.07 0.09 0.06 0.04 0.09 0.07 RD5@1 1.0 0.41 0.46 0.61 0.5 0.52 0.43 0.3 0.52 0.77 0.39 0.59 RR5@1 0.02 0.07 0.04 0.08 0.06 0.07 0.05 0.05 0.07 0.12 0.05 0.11 SynonymSubstitution RP5@1 0.01 0.04 0.02 0.04 0.05 0.04 0.04 0.05 0.05 0.04 0.08 0.07 RD5@1 0.5 0.59 0.77 0.61 0.55 0.67 0.67 0.55 0.62 0.73 0.43 0.56 RR5@1 0.02 0.07 0.06 0.09 0.07 0.09 0.09 0.07 0.08 0.12 0.07 0.1 WhitespacePerturbation RP5@1 0.0 0.05 0.02 0.04 0.05 0.04 0.06 0.06 0.05 0.04 0.07 0.05 RD5@1 1.0 0.53 0.69 0.67 0.55 0.67 0.52 0.5 0.62 0.77 0.52 0.67 RR5@1 0.02 0.08 0.05 0.09 0.07 0.1 0.07 0.06 0.08 0.12 0.07 0.12 SynonymInsertion RP5@1 0.01 0.05 0.02 0.03 0.06 0.05 0.09 0.07 0.05 0.07 0.09 0.03 RD5@1 0.75 0.47 0.69 0.72 0.5 0.62 0.33 0.4 0.62 0.54 0.39 0.81 RR5@1 0.02 0.06 0.07 0.09 0.06 0.08 0.05 0.05 0.08 0.1 0.07 0.15
Table 14: Performance of code generation models against function name perturbations. (HumanEval-X and worst case passes)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal RP5@1 0.02 0.1 0.09 0.11 0.12 0.13 0.13 0.13 0.13 0.16 0.15 0.16
FuncRenameSynonymSub RP5@1 0.01 0.03 0.04 0.04 0.07 0.07 0.08 0.07 0.04 0.05 0.07 0.09
RD5@1 0.75 0.71 0.54 0.67 0.45 0.48 0.38 0.45 0.67 0.69 0.48 0.48
RR5@1 0.02 0.09 0.05 0.07 0.07 0.06 0.05 0.07 0.09 0.11 0.07 0.09
FuncRenameInflectionalVariation RP5@1 0.0 0.03 0.04 0.04 0.07 0.07 0.09 0.07 0.07 0.03 0.09 0.07
RD5@1 1.0 0.71 0.54 0.67 0.4 0.43 0.29 0.45 0.43 0.81 0.39 0.59
RR5@1 0.02 0.07 0.04 0.07 0.06 0.08 0.05 0.05 0.07 0.13 0.05 0.1
FuncRenameChangeChar RP5@1 0.0 0.02 0.05 0.05 0.07 0.06 0.06 0.04 0.07 0.05 0.07 0.07
RD5@1 1.0 0.82 0.31 0.56 0.45 0.52 0.52 0.65 0.43 0.69 0.48 0.56
RR5@1 0.02 0.1 0.05 0.07 0.05 0.08 0.07 0.08 0.07 0.13 0.08 0.12
FuncRenameButterFinger RP5@1 0.0 0.05 0.03 0.03 0.04 0.07 0.07 0.08 0.07 0.04 0.07 0.07
RD5@1 1.0 0.47 0.62 0.72 0.65 0.48 0.43 0.35 0.48 0.73 0.48 0.56
RR5@1 0.02 0.09 0.05 0.08 0.08 0.07 0.05 0.04 0.07 0.12 0.07 0.1
FuncRenameSwapChar RP5@1 0.01 0.05 0.04 0.05 0.06 0.07 0.07 0.07 0.07 0.05 0.09 0.07
RD5@1 0.75 0.53 0.46 0.56 0.5 0.48 0.43 0.4 0.43 0.69 0.35 0.56
RR5@1 0.02 0.07 0.05 0.07 0.06 0.07 0.05 0.06 0.07 0.11 0.06 0.12
FuncRenameSnake/CamelCase RP5@1 0.01 0.06 0.04 0.04 0.07 0.07 0.08 0.09 0.07 0.07 0.09 0.07
RD5@1 0.75 0.41 0.46 0.67 0.4 0.43 0.38 0.3 0.48 0.54 0.39 0.59
RR5@1 0.02 0.07 0.05 0.09 0.07 0.05 0.05 0.05 0.06 0.09 0.07 0.11
Table 15: Performance of code generation models against syntax perturbations. (HumanEval-X and worst case passes)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS Nominal RP5@1 0.34 0.1 0.37 0.46 0.16 0.43 0.57 0.11 0.41 0.55 0.2 0.44 OperandSwap RP5@1 0.29 0.04 0.29 0.32 0.08 0.33 0.37 0.1 0.21 0.35 0.1 0.29 RD5@1 0.11 0.65 0.25 0.27 0.54 0.25 0.34 0.24 0.45 0.35 0.53 0.35 RR5@1 0.2 0.08 0.11 0.18 0.13 0.11 0.2 0.08 0.18 0.19 0.14 0.19 ForWhileTransformer RP5@1 0.26 0.04 0.23 0.3 0.05 0.27 0.37 0.1 0.22 0.32 0.13 0.25 RD5@1 0.22 0.65 0.41 0.33 0.71 0.39 0.34 0.19 0.44 0.4 0.42 0.43 RR5@1 0.21 0.08 0.17 0.21 0.15 0.17 0.23 0.09 0.2 0.21 0.12 0.23 VarRenamerNaive RP5@1 0.27 0.02 0.27 0.34 0.06 0.29 0.38 0.07 0.24 0.32 0.1 0.27 RD5@1 0.19 0.76 0.3 0.25 0.64 0.33 0.32 0.43 0.38 0.41 0.56 0.39 RR5@1 0.17 0.08 0.12 0.21 0.13 0.15 0.2 0.07 0.16 0.22 0.13 0.21 DeadCodeInserter RP5@1 0.04 0.01 0.08 0.05 0.02 0.1 0.07 0.02 0.06 0.07 0.04 0.09 RD5@1 0.89 0.88 0.79 0.89 0.89 0.78 0.87 0.81 0.84 0.88 0.81 0.81 RR5@1 0.29 0.09 0.3 0.4 0.15 0.34 0.49 0.12 0.33 0.47 0.19 0.38 VarRenamerRN RP5@1 0.26 0.03 0.26 0.31 0.04 0.3 0.38 0.07 0.24 0.31 0.09 0.26 RD5@1 0.22 0.71 0.32 0.3 0.75 0.32 0.33 0.48 0.39 0.42 0.61 0.42 RR5@1 0.2 0.07 0.12 0.22 0.15 0.14 0.22 0.09 0.15 0.23 0.13 0.22 VarRenamerCB RP5@1 0.27 0.03 0.27 0.33 0.07 0.32 0.39 0.05 0.22 0.37 0.1 0.26 RD5@1 0.17 0.71 0.29 0.26 0.57 0.28 0.3 0.57 0.44 0.32 0.56 0.42 RR5@1 0.21 0.09 0.11 0.2 0.15 0.12 0.21 0.07 0.17 0.18 0.13 0.23
Table 16: Performance of code generation models against format perturbations.
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal RP5@1 0.34 0.1 0.37 0.46 0.16 0.43 0.57 0.11 0.41 0.55 0.2 0.44
split_lines RP5@1 0.3 0.03 0.28 0.37 0.07 0.32 0.47 0.1 0.27 0.43 0.1 0.26
RD5@1 0.07 0.71 0.27 0.16 0.57 0.26 0.16 0.19 0.31 0.2 0.53 0.4
RR5@1 0.12 0.09 0.13 0.09 0.1 0.13 0.1 0.05 0.16 0.11 0.12 0.19
new_line_afterdoc RP5@1 0.25 0.04 0.2 0.38 0.09 0.37 0.5 0.12 0.3 0.46 0.14 0.21
RD5@1 0.24 0.65 0.48 0.14 0.5 0.17 0.11 0.1 0.22 0.14 0.36 0.51
RR5@1 0.12 0.07 0.18 0.11 0.1 0.07 0.09 0.02 0.12 0.1 0.08 0.25
new_line_aftercode RP5@1 0.2 0.04 0.18 0.36 0.07 0.35 0.42 0.09 0.16 0.45 0.12 0.18
RD5@1 0.39 0.59 0.52 0.19 0.61 0.19 0.25 0.29 0.59 0.16 0.47 0.58
RR5@1 0.16 0.07 0.21 0.16 0.12 0.11 0.16 0.04 0.26 0.11 0.1 0.29
tab_indent RP5@1 0.12 0.04 0.35 0.08 0.09 0.41 0.49 0.1 0.34 0.43 0.14 0.37
RD5@1 0.63 0.59 0.08 0.82 0.46 0.07 0.12 0.24 0.14 0.19 0.36 0.15
RR5@1 0.21 0.06 0.04 0.37 0.08 0.05 0.08 0.03 0.09 0.1 0.09 0.14
doc2comments RP5@1 0.1 0.03 0.39 0.2 0.06 0.02 0.35 0.1 0.38 0.4 0.12 0.41
RD5@1 0.69 0.71 -0.02 0.56 0.64 0.94 0.38 0.24 0.03 0.25 0.47 0.06
RR5@1 0.23 0.07 0.04 0.25 0.13 0.41 0.21 0.04 0.09 0.15 0.1 0.07
new_lines RP5@1 0.24 0.04 0.14 0.32 0.07 0.32 0.46 0.09 0.18 0.36 0.12 0.15
RD5@1 0.28 0.59 0.63 0.29 0.61 0.26 0.18 0.29 0.55 0.33 0.47 0.65
RR5@1 0.13 0.07 0.24 0.15 0.13 0.12 0.12 0.04 0.26 0.19 0.1 0.31
Table 17: P Values Using Fisher’s Exact Test
HumanEval-X Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Java CPP JS Java CPP JS Java CPP JS Java CPP JS
DocString 0.25 0.08 0.03 0.01 0.08 0.02 0.09 0.06 0.02 0.0 0.05 0.0
Syntax 0.04 0.01 0.0 0.0 0.0 0.0 0.0 0.11 0.0 0.0 0.0 0.0
Format 0.0 0.03 0.0 0.0 0.02 0.01 0.0 0.58 0.0 0.0 0.03 0.0
Function 0.25 0.08 0.1 0.02 0.08 0.06 0.14 0.06 0.06 0.0 0.08 0.01
6 Threats to Validity
Internal Validity: Threats to internal validity indicate the internal factors that may affect the result. For this work, generating the same codes from the same prompt each time by the code generation model can introduce biases into the result. To eliminate this, we generate 5 perturbed datasets from the original dataset, and all are used to mitigate the randomness of the code generation model. We run the whole process for all 28 types of perturbed datasets, 5 times each. Besides, the dataset we used is a public benchmark, and we have no control over it, which could impact the result.
External Validity: Threats to external validity indicate outside effects other than the codes and used dataset. In this work as we use three code generation models and the human-eval-x dataset, which consists of datasets for multiple languages. the dataset and model, which could introduce bias for a limited number of models. However, we prepared our approach in a generic way that could be used for other languages and other datasets also to mitigate this threat.
7 Conclusion and Future Work
This paper addresses the demand for a comparative study on the robustness of code generation models across multiple popular programming languages, including Java, C++, and JavaScript. Building upon the foundation of ReCode, which primarily focuses on Python, we extend our work to encompass these languages. Through our research, we adapt 28 distinct types of perturbations from ReCode within four scopes of the prompt. Our investigation yields intriguing findings, which we present in this paper. Notably, we observe that code generation models exhibit varying behaviors across different programming languages and in response to different perturbations. The inherent characteristics of the programming languages have a notable impact on the models’ performance. However, conducting such black box experiments comes with significant computational costs. Therefore, as a future endeavor, we aim to explore white box strategies to enhance the robustness analysis of code generation models.